OpenStudy (anonymous):

1.)G(x)=√(4-x) a.G(4-x) b.G(4-x^2 ) c.G(4-x^4 ) d.G(4x-x^2 ) e.G(-x^4-4x^2)

OpenStudy (anonymous):

What's the question? .... Which of those options is G(x) = sqrt(4-x) equivalent to/

OpenStudy (anonymous):

Evaluate the function at the indicated values

OpenStudy (anonymous):

oh! derp. That makes sense. Are you having trouble with one in particular?

OpenStudy (anonymous):

with all.. :/

OpenStudy (anonymous):

:) none worries. Evaluating a function at a location just means substituting that location wherever x is in the original equation. So for the first one you take your original equation \[ G_{(x)}= \sqrt{4-(x)}\] and then replace that x with the expression (4-x), \[ G_{(4-x)}= \sqrt{4-(4-x)}\] Then simplify. You try #2 :)

OpenStudy (anonymous):

wait.. how to simplify?

OpenStudy (anonymous):

\[ G_{(4-x)}= \sqrt{4-4+x)}\] \[ G_{(4-x)}= \sqrt{x}\]

OpenStudy (anonymous):

let me try number 2. :)

OpenStudy (anonymous):